# Models in Long-Context Interaction

### v1

@@ Abstract
Large language models (LLMs) make significant progress in Emotional Intelligence (EI) and long-context understanding. However, existing benchmarks tend to overlook certain aspects of EI in long-context scenarios, especially under realistic, practical settings where interactions are lengthy, diverse, and often noisy. To move towards such realistic settings, we present LongEmotion, a benchmark specifically designed for long-context EI tasks. It covers a diverse set of tasks, including Emotion Classification, Emotion Detection, Emotion QA, Emotion Conversation, Emotion Summary, and Emotion Expression. On average, the input length for these tasks reaches 8,777 tokens, with long-form generation required for Emotion Expression. To enhance performance under realistic constraints, we incorporate Retrieval-Augmented Generation (RAG) and Collaborative Emotional Modeling (CoEM), and compare them with standard prompt-based methods. Unlike conventional approaches, our RAG method leverages both the conversation context and the large language model itself as retrieval sources, avoiding reliance on external knowledge bases. The CoEM method further improves performance by decomposing the task into five stages, integrating both retrieval augmentation and limited knowledge injection. Experimental results show that both RAG and CoEM consistently enhance EI-related performance across most long-context tasks, advancing LLMs toward more practical and real-world EI applications. Furthermore, we conducted a comparative case study experiment on the GPT series to demonstrate the differences among various models in terms of EI.
[…]
Codehttps://github.com/LongEmotion/LongEmotion
[…]
Projecthttps://longemotion.github.io/
[…]
Emailweichuliu1023@gmail.com
[…]
Emailjunexiong@connect.hku.hk
@@ 1 Introduction
Large Language Models (LLMs) are increasingly adopted in the domain of Emotional Intelligence (EI) (Wang et al. 2023; Sabour et al. 2024). For instance, the EmoBench (Sabour et al. 2024) highlights the necessity for robust, psychological-theory-grounded evaluation across both emotional understanding and generation, demonstrating that current LLMs lag behind human-level performance. By leveraging their advanced language understanding and generation capabilities, LLMs become valuable tools for facilitating emotional expression (Ishikawa and Yoshino 2025; Lu et al. 2025), with recent work showing their capacity to simulate specified emotional states in accordance with established models such as Russell’s Circumplex (Russell 1980, 2003). LLMs are increasingly serving in roles ranging from mental health assistants (Guo et al. 2024; Malgaroli et al. 2025; Fu et al. 2024) to everyday conversational companions (Fu et al. 2024; Duan et al. 2024; Zhang et al. 2025). This growing integration into emotionally sensitive domains places greater demand on LLMs to maintain emotional coherence over time — not only to understand but also to remember, adapt, and respond empathetically in prolonged interactions (Zhong et al. 2024).
In particular, during long-context interactions (Maharana et al. 2024), LLMs are expected to recognize emotional cues embedded across temporally dispersed user inputs and to deliver nuanced, empathetic responses that reflect continuity in emotional expression. As such, users increasingly turn to LLM-based chatbots for both knowledge and emotional support in dynamic, evolving conversations.
@@
Figure 1: (a) Token distributions across tasks. For Emotion Expression, the sequence length refers to the average length of model-generated outputs, whereas for the other tasks, it corresponds to the average length of input contexts. (b) Distribution of sample counts across the six tasks, illustrating the overall composition of the dataset.
However, existing benchmarks (Sabour et al. 2024; Maharana et al. 2024; Paech 2023; Hu et al. 2025) for assessing the EI of LLMs lack the contextual and temporal depth required to evaluate EI in long-context settings. While these efforts lay a strong foundation, they fall short in capturing the nuanced, dynamic nature of emotional communication as it occurs in real-world long-contexts (Xiong et al. 2025, 2024). More specifically: i) In most current benchmarks, emotion recognition tasks rely on input texts that are short, explicit, and often directly labeled with clear emotional cues. This simplifies the task and does not reflect natural conversations, where emotional content is frequently subtle, embedded across multiple turns, and obscured by irrelevant or noisy information. ii) Existing generation-based benchmarks largely focus on short, multi-turn dialogues with a limited number of conversational turns. As a result, they fail to challenge LLMs to maintain emotional coherence over extended interactions. iii) Moreover, an equally important but underexplored aspect is the LLM’s ability to generate its own emotional expressions in long-form outputs, not just recognize or respond to others—an area still lacking robust evaluation. iv) The capacity of LLMs to leverage internalized emotional knowledge—such as theoretical emotion models, social-emotional reasoning, or culturally grounded affective norms—is crucial to demonstrating higher-order EI. Yet, current evaluations rarely include tasks that assess how models apply such knowledge across longer contexts or evolving emotional trajectories. Given these limitations, can existing benchmarks truly capture the core dimensions of Emotional Intelligence—such as perception, expression, and regulation—within realistic long-context settings?
@@
To bridge the gap between realistic scenarios and long-context evaluation, we introduce LongEmotion, a benchmark designed to mirror real-world conversational dynamics when assessing LLMs’ EI over long-context interactions. LongEmotion comprises six complementary tasks. Four tasks—Emotion Classification, Emotion Detection, Emotion Conversation, and Emotion Expression—measure a model’s ability to recognize and generate emotional content when context spans many dialogue turns or involves complex scenarios. Two knowledge-intensive tasks—Emotion QA and Emotion Summary—probe how effectively a model leverages and applies its internalized emotional knowledge in authentic scenarios. Figure 1 depicts the dataset’s distribution, and a high‐level overview appears in Figure 2.
@@
To handle these realistic settings, we develop a Retrieval‐Augmented Generation (RAG) approach as well as a novel multi‐agent emotional modeling framework called Collaborative Emotional Modeling (CoEM). Unlike standard RAG systems that pull from static, external corpora, our method treats the conversation history itself as a dynamic vector store to capture aspect-level sentiment terms. To further enhance long-context emotional understanding, we introduce CoEM, where the context is divided into coherent chunks, roughly ranked by relevance, and then processed by multiple collaborating agents (e.g., an auxiliary GPT-4o instance (OpenAI 2024b)). After a second‐stage re‐ranking, these agents collectively generate an emotional “ensemble” response. This pipeline not only reflects the unpredictability and noise of real‐world dialogue but also emphasizes how emotionally salient information can be continuously extracted, re‐contextualized, and articulated over long-context interaction. Our contributions are summarized as:
@@
  • We present LongEmotion, a long-context EI benchmark with six diverse tasks targeting recognition, generation, and knowledge application.
@@
  • We propose RAG and CoEM frameworks to enhance performance by retrieving and enriching contextually relevant information.
@@
  • We perform extensive experiments across all settings, offering detailed analyses of LLMs’ EI in long-context scenarios.
@@ 2 Related Work
Figure 2: An illustrative overview of the LongEmotion dataset. To comprehensively evaluate the Emotional Intelligence of LLMs in long-context interaction, we design six tasks: Emotion Classification, Emotion Detection, Emotion QA, Emotion Conversation, Emotion Summary, and Emotion Expression.
@@
Many benchmarks are developed to assess LLMs’ Emotional Intelligence (EI). EmoBench (Sabour et al. 2024; Hu et al. 2025) draws on psychological theories to evaluate both emotional understanding and application across 400 English–Chinese handcrafted questions, exposing significant gaps between model and human EI levels. EQ‑Bench (Paech 2023) measures LLMs’ ability to rate emotional intensity in dialogues through 60 English queries, showing strong correlation with multi-domain reasoning benchmarks. More recently, EmotionQueen (Chen et al. 2024b) offers a specialized benchmark for empathy, requiring LLMs to recognize key events, implicit emotions, and generate empathetic responses. Despite their strengths, all of these focus on short or synthetic interactions and lack the long contextual depth critical for assessing EI in extended conversational or narrative settings.
Long-Context Understanding.
LLMs make strides in processing long documents, yet robust evaluation remains an open challenge. LongBench (Bai et al. 2023) introduces a bilingual, multi-task benchmark covering QA, summarization, and code tasks with average context lengths over 6,000 words, revealing that even state-of-the-art models struggle with extended inputs. Complementing this, LooGLE (Li et al. 2023) evaluates long-context reasoning using realistic documents exceeding 24k tokens, uncovering dependencies that span across distant spans. For extreme-length evaluation, XL2Bench (Ni et al. 2024) includes tasks on fiction, law, and scientific papers with inputs up to 100k+ words—yet LLMs still fall short in handling long-range dependencies.
Beyond these, RULER (Chen et al. 2023) focuses on complex reasoning chains in long-form texts via fine-grained question types and inter-paragraph dependencies, providing a valuable diagnostic lens into model reasoning depth. InfiniteBench (Sun, Gao et al. 2024), meanwhile, evaluates LLMs’ abilities on open-ended, unbounded contexts with theoretically unlimited input lengths, highlighting model degradation as input exceeds trained context windows.
Survey work such as Liu et al. (2025) offers a broad overview of long-context modeling and evaluation paradigms but emphasizes that most benchmarks primarily target information retrieval or general comprehension—not emotional intelligence or affective computing.
@@ 3 LongEmotion: Construction and Task
Building on LongEmotion, we evaluate models’ EI capabilities using three prompt-based methods: Base, RAG, and CoEM. The statistical overview of LongEmotion dataset can be found in Table 1. Appendix F provides a detailed explanation of metrics used in tasks where LLMs act as evaluators.
3.1 Emotion Classification
Task Design.
This task requires the model to identify the emotional category of a target entity within long-context texts that contain lengthy spans of context-independent noise (Kamradt 2023). Model performance is evaluated based on the consistency between the predicted label and the ground truth.
Data Construction.
We embed short emotional excerpts from Emobench into passages drawn from BookCorpus (Zhu et al. 2015). The construction involves: i) random insertion of emotional snippets, and ii) manual adjustments to ensure syntactic and contextual coherence. Proper nouns are modified to avoid identity overlap, while preserving the intended emotional tone.
3.2 Emotion Detection
Task Design.
@@ 3.2 Emotion Detection
Dataset Construction.
Construction follows two steps: i) grouping texts from the Covid-worry dataset by emotion label, and ii) inserting a mismatched segment into each group to form challenging contrast sets.
    Task ID Construction Method Metric Avg len Count
    Emotion Classification EC Segment Insertion & Manual Adjustment Accuracy 16691 200
    Emotion Detection ED Reorganization of Existing Dataset Accuracy 4106 136
    Emotion QA QA Human Annotation F1 11207 120
    Emotion Conversation MC Model Expansion & Human Annotation LLM as Judge 4856 100
    Emotion Summary ES Model Expansion & Human Annotation LLM as Judge 3129 150
    Emotion Expression $E$ Reorganization of Existing Dataset LLM as Judge 8546* 428
Table 1: A statistical overview of the LongEmotion dataset. ID represents the abbreviations of each task. Among them, EC, ED, QA, MC, and ES are long text input tasks, where Avg len refers to the average context length of the entries. $E$ is a long text generation task, with Avg len indicating models’ average generation length, and an asterisk (*) placed at the top right corner for special notation. LLM as Judge indicates scoring performed by GPT-4o.
3.3 Emotion Conversation
Task Design.
In our four-stage long-context dialogue dataset, we select the quartile, half, and three-quarter points of each stage as evaluation checkpoints to assess the model’s EI capabilities. At these checkpoints, the model is required to act as a psychological counselor and provide empathetic and contextually appropriate emotional support to the patient. Following the generation of model responses, we conduct a comprehensive evaluation using stage-specific metrics meticulously designed from the perspective of professional psychological counseling. The scoring is performed by GPT-4o, which serves as the evaluator to ensure consistency and scalability in assessment. To better highlight the advantages of our CoEM in long-context scenarios, we only apply the RAG and CoEM methods in the fourth stage of the dialogue. We evaluate the performance of all models across all stages under the Base setting.
Dataset Construction.
Based on CPsyCoun (Zhang et al. 2024), we construct 100 emotionally rich dialogues by expanding seed prompts into four functional stages: i) Reception and Inquiry, ii) Diagnostic, iii) Consultation, and iv) Consolidation and Ending. Each stage reflects key elements in therapeutic progression.
To assess model performance across stages, we introduce 12 specialized metrics informed by five major therapeutic frameworks: Cognitive Behavioral Therapy (CBT) (Beck 2021), Acceptance and Commitment Therapy (ACT) (Waltz and Hayes 2010), Humanistic Therapy (Elliott 2002), Existential Therapy (May 1994), and Satir Family Therapy (Rebner 1972). The complete definitions of these 12 metrics are in Appendix F. To ensure high dataset quality, we conduct a quality assessment using two parallel evaluation protocols: i) manual scoring by psychology experts, and ii) automated assessment using GPT-4o. Table 2 reports the average quality performance by stage, where the pearson correlation coefficient between LLMs and human scores reaches 0.934 (p = 0.066), indicating a strong alignment between automated and manual evaluations. We use inter-annotator agreement (Fleiss 1971) to measure the consistency among annotators, as shown in Appendix B. For the qualifications of psychology annotators, please refer to Appendix A.
Stage Evaluator
GPT-4o Annotator
Reception and Inquiry Stage 4.36 3.89
Diagnostic Stage 4.06 3.86
Consultation Stage 3.79 3.65
Consolidation and Ending Stage 4.56 3.96
Table 2: Quality Evaluation on Emotion Conversation.
3.4 Emotion QA
Task Design.

### v2

@@ Abstract
Large language models (LLMs) make significant progress in Emotional Intelligence (EI) and long-context modeling. However, existing benchmarks often overlook the fact that emotional information processing unfolds as a continuous long-context process. To address the absence of multidimensional EI evaluation in long-context inference and explore model performance under more challenging conditions, we present LongEmotion, a benchmark that encompasses a diverse suite of tasks targeting the assessment of models’ capabilities in Emotion Recognition, Knowledge Application, and Empathetic Generation, with an average context length of 15,341 tokens. To enhance performance under realistic constraints, we introduce the Collaborative Emotional Modeling (CoEM) framework, which integrates Retrieval-Augmented Generation (RAG) and multi-agent collaboration to improve models’ EI in long-context scenarios. We conduct a detailed analysis of various models in long-context settings, investigating how reasoning mode activation, RAG-based retrieval strategies, and context-length adaptability influence their EI performance.
LongEmotion: Measuring Emotional Intelligence of Large Language
Models in Long-Context Interaction
Weichu Liu1, Jing Xiong2, Yuxuan Hu3, Zixuan Li4, Minghuan Tan5, Ningning Mao6, Hui Shen2,7, Wendong Xu2, Chaofan Tao2, Min Yang5, Chengming Li1, Lingpeng Kong2, Ngai Wong2 1Shenzhen MSU-BIT University, 2The University of Hong Kong, 3City University of Hong Kong, 4Institute of Automation, Chinese Academy of Sciences, 5Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences, 6Beijing Normal University, 7University of Michigan, Ann Arbor  Github: https://longemotion.github.io/[Uncaptioned image] Huggingface: LongEmotion
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly adopted in the domain of Emotional Intelligence (EI) (Wang et al., 2023). By leveraging their advanced language understanding and generation capabilities, LLMs become valuable tools for facilitating emotional expression (Ishikawa and Yoshino, 2025; Lu et al., 2025), with recent work showing their capacity to simulate specified emotional states in accordance with established models such as Russell’s Circumplex (Russell, 1980, 2003). LLMs are increasingly serving in roles ranging from mental health assistants (Guo et al., 2024; Malgaroli et al., 2025; Fu et al., 2024) to everyday conversational companions (Fu et al., 2024; Duan et al., 2024; Zhang et al., 2025). This growing integration into emotionally sensitive domains places greater demand on LLMs to maintain emotional coherence over time — not only to understand but also to remember, adapt, and respond empathetically in prolonged inference (Zhong et al., 2024).
Although existing benchmarks make considerable progress in measuring the EI of LLMs (Sabour et al., 2024; Huang et al., 2024), current evaluation still suffers from the following limitations: (i) As articulated by Affective Information Processing Theory (Lang and Cuthbert, 1984), humans continuously receive, process, organize, and respond to emotional information, which can manifest unique patterns of emotional intelligence within a long-context setting. Existing studies often overlook the gap between idealized conditions and real-world scenarios: in realistic settings, the processing of emotional information is a continuous and enduring process. To bridge this gap, models should be evaluated on their EI in long context, which can be further decomposed into three key abilities: accurate emotion recognition, appropriate knowledge application, and affectively empathetic expression in long-context inference. (ii) Current research predominantly focuses on measuring a single aspect of the model’s capabilities, such as classification, expression, etc. According to the Mayer-Salovey-Caruso Emotional Intelligence Test (MSCEIT) (Mayer et al., 2002), an individual’s EI encompasses multiple dimensions. Assessing only one specific capability is insufficient to fully represent the model’s EI. (iii) The emotional behavior of recent state-of-the-art techniques in long-context scenarios remains unexplored, especially for reasoning models with think mode, RAG-based agent methods, and other emerging approaches.
@@
Figure 1: (a) Sequence length denotes average model output length for Emotion Expression, and average input context length for other tasks. (b) Distribution of sample counts across the six tasks, illustrating the overall composition of the dataset.
@@
To bridge realistic scenarios and long-context evaluation, we introduce LongEmotion, a benchmark designed to comprehensively evaluate the EI of LLMs in long-context inference. LongEmotion comprises six complementary tasks. Two Emotion Recognition tasks, Emotion Classification and Emotion Detection, measure the model’s reasoning ability when key emotional information is located in noisy, long-context scenarios; two Empathetic Generation tasks, Emotion Conversation and Emotion Expression, evaluate the model’s empathy and expression abilities in the context of expansive multi-turn conversations or self-narratives; two Knowledge Application tasks, Emotion QA and Emotion Summary, probe how effectively the model leverages and applies emotional knowledge in authentic scenarios. Figure 1 depicts the dataset’s distribution.
@@
To handle these realistic settings, we develop a Retrieval‐Augmented Generation (RAG) approach as well as a novel multi‐agent emotional modeling framework called Collaborative Emotional Modeling (CoEM). Unlike standard RAG systems that pull from static, external corpora, our method treats the conversation history itself as a dynamic vector store to capture aspect-level sentiment terms. To further enhance EI in long context, we introduce CoEM, where the context is divided into coherent chunks, initially ranked by relevance, and then processed by multiple collaborating agents (e.g., an auxiliary GPT-4o instance (OpenAI, 2024b)). After a second‐stage re‐ranking, these agents collectively generate an emotional “ensemble” response. This architecture captures the uncertainty and fluidity of real-world dialogue, allowing emotionally salient information to be continuously extracted, re-contextualized, and articulated. To further investigate the applicability of RAG techniques to long-context-based emotional tasks, we also adapt Self-RAG (Asai et al., 2024) and Search-o1 (Li et al., 2025) methods to LongEmotion by replacing their retrieval corpus with conversational context, thereby exploring broader possibilities of RAG in the domain of Emotional Intelligence. Our contributions are summarized as:
@@
  • We present LongEmotion, a long-context EI benchmark with six diverse tasks targeting recognition, generation, and psychological knowledge application.
@@
  • We propose CoEM framework to enhance performance by retrieving and enriching contextually relevant information.
@@
  • We perform extensive experiments across all settings and comprehensive case study, offering detailed analyses of LLMs’ EI in long-context scenarios.
@@ 2 Related Work
Figure 2: An illustrative overview of the LongEmotion dataset. To comprehensively evaluate the EI of LLMs in long-context interaction, we design six tasks: Emotion Classification, Emotion Detection, Emotion QA, Emotion Conversation, Emotion Summary, and Emotion Expression.
@@
Many benchmarks are developed to assess LLMs’ Emotional Intelligence (EI). Emobench (Sabour et al., 2024) draws on psychological theories to evaluate both emotional understanding and application across 400 English–Chinese handcrafted questions, exposing significant gaps between model and human EI levels. EQ‑Bench (Paech, 2023) measures LLMs’ ability to rate emotional intensity in dialogues through 60 English queries, showing strong correlation with multi-domain reasoning benchmarks. More recently, EmotionQueen (Chen et al., 2024b) offers a specialized benchmark for empathy, requiring LLMs to recognize key events, implicit emotions, and generate empathetic responses. Despite their strengths, all of these focus on short or synthetic interactions and lack the long contextual depth critical for assessing EI in extended conversational or narrative settings.
Long-Context Modeling.
LLMs make strides in processing long documents, yet robust evaluation remains an open challenge. LongBench (Bai et al., 2023) introduces a bilingual, multi-task benchmark covering QA, summarization, and code tasks with average context lengths over 6,000 words, revealing that even state-of-the-art models struggle with extended inputs. Complementing this, LooGLE (Li et al., 2023) evaluates long-context reasoning using realistic documents exceeding 24k tokens, uncovering dependencies that span across distant spans. For extreme-length evaluation, XL2Bench (Ni et al., 2024) includes tasks on fiction, law, and scientific papers with inputs up to 100k+ words—yet LLMs still fall short in handling long-range dependencies. Beyond these, RULER (Chen et al., 2023) focuses on complex reasoning chains in long-form texts via fine-grained question types and inter-paragraph dependencies, providing a valuable diagnostic lens into model reasoning depth. InfiniteBench (Sun et al., 2024), meanwhile, evaluates LLMs’ abilities on open-ended, unbounded contexts with theoretically unlimited input lengths, highlighting model degradation as input exceeds trained context windows. Survey work such as Liu et al. (2025) offers a broad overview of long-context modeling and evaluation paradigms but emphasizes that most benchmarks primarily target information retrieval or general comprehension—not emotional intelligence or affective computing.
@@ 3 LongEmotion: Construction and Task
A visual overview is shown in Figure 2. Appendix D provides a detailed explanation of metrics used in tasks where LLMs act as evaluators. We summarize the advantages of LongEmotion in enhancing LLMs’ EI in Appendix B.3
Task ID Source Construction Metric Avg len Count
[…]
Emotion Classification EC Emobench, FinEntity Segment Insertion Accuracy 30139 400
Emotion Detection ED Covid-worry Reorganization Accuracy 4106 136
[…]
Emotion QA QA Literature Human Annotation F1 11207 120
Emotion Summary ES CPsycoun Human Annotation LLM as Judge 15341 150
[…]
Emotion Conversation MC CPsycoun Expansion LLM as Judge 4856 100
Emotion Expression $E$ EmotionBench Reorganization LLM as Judge 8546* 428
Table 1: A statistical overview of the LongEmotion dataset. ID denotes task abbreviations. EC, ED, QA, MC, and ES involve long-text input, with Avg len showing average context length. $E$ is a long-text generation task—Avg len here refers to average output length (marked with *).
3.1 Task Design
Emotion Classification.
This task requires the model to identify the emotional category of a target entity within long-context texts that contain lengthy spans of context-independent noise (Kamradt, 2023). Model performance is evaluated by its accuracy against the ground truth.
Emotion Detection.
@@ 3.1 Task Design
Emotion QA.
@@
Emotion Summary.
In this task, the model is required to summarize the following aspects from long-context psychological pathology reports: (i) causes, (ii) symptoms, (iii) treatment process, (iv) illness characteristics, and (v) treatment effects. After generating the model’s response, we employ GPT-4o to evaluate its factual consistency, completeness, and clarity with respect to the reference answer. These three evaluation criteria are validated in CPsyExam (Zhao et al., 2024).
Emotion Conversation.
In our four-stage long-context counseling dialogue dataset, we select the quartile, half, and three-quarter points of each stage as evaluation checkpoints to assess the model’s EI capabilities. We introduce 12 specialized metrics informed by five major therapeutic frameworks: Cognitive Behavioral Therapy (CBT) (Beck, 2021), Acceptance and Commitment Therapy (ACT) (Waltz and Hayes, 2010), Humanistic Therapy (Elliott, 2002), Existential Therapy (May, 1994), and Satir Family Therapy (Rebner, 1972), which can be seen in Appendix D. The scoring is performed by GPT-4o, which serves as the evaluator to ensure consistency and scalability.
Emotion Expression.
3.2 Data Construction
The statistical overview of LongEmotion dataset can be found in Table 1. EC and ED tasks focus on evaluating the model’s ability in emotional recognition. QA and ES tasks emphasize the model’s capability to apply knowledge within long-context scenarios. MC and EE tasks aim to measure the model’s generative ability.
Reorganization from Existing Datasets.
In Emotion Classification, we embed short excerpts from Emobench (Sabour et al., 2024) and FinEntity (Tang et al., 2023) into BookCorpus passages (Zhu et al., 2015), by randomly inserting snippets and manually adjusting proper nouns for coherence. In Emotion Detection, we build contrast sets by grouping texts from Covid-worry (Kleinberg et al., 2020; van der Vegt and Kleinberg, 2023) by emotion label and inserting mismatched segments. In Emotion Expression, we use situations from EmotionBench (Huang et al., 2024) to provide models with specific emotional contexts.
Expansion and Human Annotation
For Emotion Conversation, based on CPsyCoun (Zhang et al., 2024), we construct 100 emotionally rich dialogues by expanding seed prompts into four functional stages: (i) Reception and Inquiry, (ii) Diagnostic, (iii) Consultation, and (iv) Consolidation and Ending. Dataset quality is evaluated through two parallel protocols: (i) manual scoring by psychology experts and (ii) automated assessment with GPT-4o. As reported in Figure 3, the Pearson correlation between LLM and human scores reaches 0.934 ($p=0.066$), indicating a relatively high alignment. In addition, we use the same prompts and GPT model for evaluation as those employed in the quality assessment, which further validates the rationality of our LLM-as-Judge setting. Annotator qualifications are detailed in Appendix A.
Figure 3: Quality Evaluation on Emotion Conversation.
In Emotion Summary, drawing on CPsyCounR dataset, we first expand the experience and reflection section of the dataset to meet our requirements for long-context inputs. Next, psychology annotators label each sample across five standardized dimensions: (i) Causes, (ii) Symptoms, (iii) Treatment Process, (iv) Illness Characteristics, and (v) Treatment Effect. Finally, by filtering samples based on format, content richness, and precision, we select a final set of 150 samples. To further extend the dataset length while preserving the original semantic integrity, we employ DeepSeek-V3 (DeepSeek-AI, 2024) to perform structured decomposition and subsequent content augmentation. In Appendix B.3, we discuss the annotation discipline for the annotation process of Emotion Summary.
In constructing Emotion QA, the annotation pipeline is illustrated in Figure 4. The construction process on psychological literature involves: (i) expert-written questions targeting emotional understanding, (ii) refinement of reference answers for clarity and consistency with F1-based evaluation, and (iii) filtering based on model performance to exclude overly ambiguous or trivial examples. Through this series of manual annotation and selection, we finally obtain 120 high-quality pairs of psychological knowledge questions and answers.